OpenStudy (emilyjones284):

Ill give everyone who answers a medal and fan (: Abigail and Bailey wash dogs to make extra money. Abigail can wash all of the dogs in 5 hours. Bailey can wash all the dogs in 3 hours. How long will it take them to wash the dogs if they work together?

OpenStudy (anonymous):

UH UH 12!

OpenStudy (emilyjones284):

ummm a correct answer

OpenStudy (anonymous):

Am I right?

OpenStudy (anonymous):

12 is rightttt

OpenStudy (emilyjones284):

its not even one of the answer choices

OpenStudy (anonymous):

a*5=A b*3=A system :D

OpenStudy (emilyjones284):

what does the little star mean?

OpenStudy (anonymous):

It means multiply

OpenStudy (anonymous):

it means multiply lol a x 5 but i dont want you to confuse with letter x.

OpenStudy (emilyjones284):

and how do i set up the systen and solve it haha

OpenStudy (anonymous):

wait i am missing something

OpenStudy (anonymous):

a+b=A

OpenStudy (anonymous):

YEAY A FAN

jimthompson5910 (jim_thompson5910):

hint: solve for t 1/5 + 1/3 = 1/t

OpenStudy (emilyjones284):

5+3=8?

OpenStudy (anonymous):

no wait i wrong again :( noooo lol

jimthompson5910 (jim_thompson5910):

it can't be 8 hours because them working together means they get the job done faster than both 3 and 5 hours (assuming they do so efficiently and don't slow each other down)

OpenStudy (emilyjones284):

im so confused haha i dont get how to solve it

OpenStudy (anonymous):

Dang i just did a problem like this a week ago grrr

jimthompson5910 (jim_thompson5910):

1/5 + 1/3 = 1/t 15t*(1/5) + 15t*(1/3) = 15t*(1/t) ... multiply everything by the LCD 15t 15t/5 + 15t/3 = 15t/t 3t + 5t = 15 t = ??

OpenStudy (emilyjones284):

well i would say 15 over 8 but theres not a choice for that

jimthompson5910 (jim_thompson5910):

15/8 = what in decimal form

OpenStudy (emilyjones284):

there is 8 over 15, 1 over 15, or 1 and 7/8ths

OpenStudy (emilyjones284):

oops there is 1/8 too

jimthompson5910 (jim_thompson5910):

15/8 = (8 + 7)/8 15/8 = 8/8 + 7/8 15/8 = 1 + 7/8

jimthompson5910 (jim_thompson5910):

so looks like they want the answer as a mixed number instead as an improper fraction

OpenStudy (emilyjones284):

ohhh okay oops i got it thanks (:

OpenStudy (anonymous):

a*5=A means abigail does all dogs in 5 hours b*3=A bailey does all dogs in 3 a+b= ? so plug in A/5+A/3=8/15 A

OpenStudy (anonymous):

i'd go with jim tho lol.

OpenStudy (emilyjones284):

thanks to all of you!

jimthompson5910 (jim_thompson5910):

you're on the right track timo86m yw